If in 2016 income was $17 trillion and rose to $18 trillion in 2017, and consumption was $12 trillion in 2016 and rose to $15 trillion in 2017, what is the marginal propensity to consumer (MPC)?
Answer #1

Marginal propensity to consume is the ratio of the change in consumption and change in income.

So, MPC = Increase in consumption / Increase income = (15 - 12)/(18-17) = 3/1 = 3